OpenStudy (anonymous):

Prove that any element x, in a group, that satisfies x^9 = e, and x^11 = e, where e is the group identity, must itself be the identity.

OpenStudy (anonymous):

so far I've got \[x^9 = e = x ^{11}\] and \[x ^{11} = x ^{2}x ^{9} = x ^{2}e= x ^{2}\]

OpenStudy (anonymous):

Im just having trouble coming up with a valid last step to show that x=e

OpenStudy (anonymous):

ah it was the inverse I wasn't considering, thanks! Silly last step had me furiously sipping my coffee and wondering what was missing. Much appreciated!

OpenStudy (anonymous):

Have you done any topology? I'm looking for someone to have a look at my question.

OpenStudy (anonymous):

Not yet unfortunately :(

OpenStudy (anonymous):

Alright, thanks for having a look though.

OpenStudy (anonymous):

Sorry I made a mistake earlier. This should work though \(e = e^{-5} = (x^{2})^{-5} = x^{-10}\) \(e = e\cdot e = x^{-10} \cdot x^{11} = x\)